# A COLIBRI Photometric Study of SN 2025bvm: A Normal, Slowly Declining Type Ia Supernova

### v1

@@ 1 INTRODUCTION
Type Ia supernovae (SNe Ia) constitute fundamental tools in modern astrophysics. Their exceptional and consistent peak luminosity has established them as powerful extragalactic distance indicators. In fact, observations of high-redshift SNe Ia provided the key evidence for the discovery of the accelerated expansion of the Universe (Riess1998; Perlmutter1999). This fundamental role as standard candles is based on the remarkable uniformity of their intrinsic photometric properties, allowing for the determination of cosmic distances on cosmological scales
@@
Extensive observational studies of large photometric and spectroscopic samples (e.g., Hamuy1996; Jha2006) have revealed notable diversity in the observable characteristics of SNe Ia. Consequently, their use as standard candles requires the application of empirical relations, such as the correlation between peak luminosity and light-curve shape (Phillips1993), for their calibration and classification. This diversity is a critical aspect, as it offers invaluable clues about the underlying explosion mechanisms and the nature of their progenitor systems (e.g., Burgaz2025; Rigault2025; Ginolin2025). Currently, two possible progenitor scenarios are under investigation: the first considers the case of a single-degenerate star (a white dwarf accreting material from a non-degenerate companion), and the second option corresponds to a double-degenerate system (the merger of two white dwarfs). Understanding the wide range of observed light curves is essential for imposing constraints on these explosion models. This diversity has led to the identification of distinct subclasses, including the overluminous and slowly declining SN 1991T-like objects (Filippenko1992a), the subluminous and rapidly declining SN 1991bg-like objects (Filippenko1992b), and transitional SNe Ia that lie between these two categories. In large, unbiased samples, such as that from the ZTF SNIa DR2 (Rigault2025), normal SNe Ia form the dominant group, accounting for approximately 70% of the population. Subluminous (SN 1991bg-like) and overluminous (SN 1991T-like) events comprise roughly 15% and 5–10% respectively, with a smaller fraction classified as peculiar (Burgaz2025). Crucially, this photometric diversity is intrinsically linked to spectroscopic differences. For instance, overluminous events like SN 1991T exhibit hotter spectra near maximum light, characterized by weak or absent Si II lines but prominent Fe III features. In contrast, subluminous events like SN 1991bg show cooler spectra with deep Ti II absorption troughs. While transitional events have been of particular interest due to their possible disconnect between photometric and spectroscopic properties (Branch2009), most SNe Ia fall into the ”normal” category, which are the most widely used as standard candles. In this context, high-cadence, multi-band photometric monitoring of nearby SNe Ia is essential, as it allows for the precise measurement of decline rates, color indices, and the presence of secondary features, such as the second maximum in the near-infrared, which are key diagnostics for the explosion physics and ejecta properties. A precise characterization of these parameters is fundamental for accurately classifying a Type Ia supernova within all currently known subclasses.
@@
In this paper, we present an extensive photometric study of the Type Ia supernova SN 2025bvm. The object was first detected on 2025 February 17 by the ATLAS survey (2025TNSTR.641....1T) in the host galaxy NGC 4156 and was spectroscopically classified as a Type Ia supernova on 2025 February 21 (Balcon2025). Spectra obtained approximately one week before maximum light, available on the Transient Name Server (TNS), clearly show characteristic Si II lines and the S II ”W-trough” feature. These spectral characteristics are highly consistent with those of normal Type Ia supernovae, showing a strong resemblance to archetypal events like SN 2011fe. Following its discovery and classification, we performed an intensive multi-band follow-up campaign with the COLIBRI robotic telescope at the Observatorio Astronómico Nacional in San Pedro Mártir, Mexico. The high-cadence data obtained allowed for a robust characterization of its photometric properties, which, as will be demonstrated, confirm its classification as a normal Type Ia supernova, consistent with the general population of these events.
@@
The structure of this paper is as follows. In Section 2, we describe the photometric observations and data processing. In Section 3, we present a detailed analysis of the light curves, including a robust methodology for reddening correction and the determination of intrinsic physical properties. Finally, in Section 4, we discuss the nature of SN 2025bvm in the context of normal Type Ia supernovae and summarize our conclusions.
@@ 2 Observations and Data Reduction
The photometric data were obtained using the 1.3-m COLIBRI robotic telescope, located at the Observatorio Astronómico Nacional de San Pedro Mártir (OAN-SPM), Baja California, Mexico. The observations were made as part of the San Pedro Mártir Reverberation Mapping program, which primarily monitored the active galactic nucleus (AGN) NGC 4151. As a result of the field coverage, the host galaxy of SN 2025bvm, NGC 4156, also fell within the same field of view SN 2025bvm was discovered by the Zwicky Transient Facility (ZTF; Bellm2019) on February 20, 2025, at coordinates RA = 12:10:48.75, DEC = +39:28:28.0 (J2000.0). The Colibri telescope monitored approximately 121 days, from February 23 to June 24, 2025, obtaining photometric data for 79 nights in the $\mathrm{B}$, $\mathrm{g^{\prime}}$, $\mathrm{r^{\prime}}$, and $\mathrm{i^{\prime}}$bands.
@@
All images obtained in these observations were processed using standard procedures, including bias subtraction and flat-field correction. This data reduction was performed using IRAF (tody1986) and Python packages, specifically Astropy (astropy2013; astropy2018; astropy2022). Instrumental magnitudes were extracted using differential photometry, which relies on comparing the supernova’s brightness to a set of reference stars within the same field. Photometric calibration used a set of standard stars from the supernova field, identified in Figure 1 and listed in Table 1. Among them, the primary references used were the stars BD+40 2507 (marked as number 2) and UI+40-42 (marked as number 3). The magnitudes for these standard stars in the corresponding filters were obtained from public catalogs such as the Sloan Digital Sky Survey (Abdurro2022) and SIMBAD astronomical database [^1] https://vizier.cds.unistra.fr/viz-bin/VizieR-3?-source=II/336/apass9.
@@
Figure 1 shows the full field of view, indicating the positions of supernova SN 2025bvm (7), its host galaxy NGC 4156 (6), the nearby galaxy NGC 4151 (1), and the aforementioned reference stars. A detailed log of all photometric observations is presented in Table 1.

### v2

# arXiv:2601.07745v2[astro-ph.HE] 02 Mar 2026

@@ 1 INTRODUCTION
Type Ia supernovae (SNe Ia) are fundamental tools in modern astrophysics. Their exceptional and consistent peak luminosity have established them as powerful extragalactic distance indicators. In fact, observations of high-redshift SNe Ia provided key evidence for the accelerated expansion of the Universe (Riess1998; Perlmutter1999). This fundamental role as standard candles is based on the remarkable uniformity of their intrinsic photometric properties, allowing for the determination of cosmic distances on cosmological scales
@@
Extensive observational studies of large photometric and spectroscopic samples (e.g., Hamuy1996; Jha2006) have revealed notable diversity in the observable characteristics of SNe Ia. Consequently, their use as standard candles requires the application of empirical relations, such as the correlation between peak luminosity and light-curve shape (Phillips1993), for their calibration and classification. This diversity is a critical aspect, as it offers invaluable clues about the underlying explosion mechanisms and the nature of their progenitor systems (e.g., Burgaz2025; Rigault2025; Ginolin2025). Currently, two possible progenitor scenarios are under investigation: the first considers the case of a single-degenerate star (a white dwarf accreting material from a non-degenerate companion), and the second option corresponds to a double-degenerate system (the merger of two white dwarfs). Understanding the wide range of observed light curves is essential for imposing constraints on these explosion models. This diversity has led to the identification of distinct subclasses, including the overluminous and slowly declining SN 1991T-like objects (Filippenko1992a), the subluminous and rapidly declining SN 1991bg-like objects (Filippenko1992b), and transitional SNe Ia that lie between these two categories. In large, unbiased samples, such as that from the ZTF SNIa DR2 (Rigault2025), normal SNe Ia form the dominant group, accounting for approximately 70% of the population. Subluminous (SN 1991bg-like) and overluminous (SN 1991T-like) events comprise roughly 15% and 5–10% respectively, with a smaller fraction classified as peculiar (Burgaz2025). Crucially, this photometric diversity is intrinsically linked to spectroscopic differences. For instance, overluminous events like SN 1991T exhibit hotter spectra near maximum light, characterized by weak or absent Si II lines but prominent Fe III features. In contrast, subluminous events, such as SN 1991bg, show cooler spectra with deep Ti II absorption troughs. While transitional events have been of particular interest of their possible disconnect between photometric and spectroscopic properties (Branch2009), most SNe Ia fall into the ”normal” category, which are the most widely used as standard candles. In this context, high-cadence, multi-band photometric monitoring of nearby SNe Ia is essential, as it allows for the precise measurement of decline rates, color indices, and the presence of secondary features, such as the second maximum in the near-infrared, which are key diagnostics for the explosion physics and ejecta properties. A precise characterization of these parameters is fundamental for accurately classifying a Type Ia supernova within all currently known subclasses.
@@
In this study, we present an extensive photometric study of the Type Ia supernova SN 2025bvm. The object was first detected on February 17, 2025, by the ATLAS survey (2025TNSTR.641....1T) in the host galaxy NGC 4156 and was spectroscopically classified as a Type Ia supernova on 2025 February 21 (Balcon2025). Spectra obtained approximately one week before maximum light, available on the Transient Name Server (TNS), clearly show characteristic Si II lines and the S II ”W-trough” feature. These spectral characteristics are highly consistent with those of normal Type Ia supernovae, showing a strong resemblance to archetypal events like SN 2011fe. Following its discovery and classification, we performed an intensive multi-band follow-up campaign with the COLIBRI robotic telescope at the Observatorio Astronómico Nacional in San Pedro Mártir, Mexico. The high-cadence data obtained allowed for a robust characterization of its photometric properties, which, as will be demonstrated, confirm its classification as a normal Type Ia supernova, consistent with the general population of these events.
@@
The remainder of this paper is organized as follows. In § 2, we describe the photometric observations and the data processing. In § 3, we present a detailed analysis of the light curves, including a robust methodology for reddening correction and the determination of intrinsic physical properties of the sources. Finally, in § 4, we discuss the nature of SN 2025bvm in the context of normal Type Ia supernovae and summarize our results.
@@ 2 Observations and Data Reduction
Photometric data were obtained using the 1.3-m COLIBRI robotic telescope and DDRAGO instrument (2022SPIE12182E..1SB; 2024SPIE13096E..3DL), located at the Observatorio Astronómico Nacional de San Pedro Mártir (OAN-SPM), Baja California, Mexico. The observations were made as part of the San Pedro Mártir Reverberation Mapping program, which primarily monitored the active galactic nucleus (AGN) NGC 4151. As a result of the field coverage, the host galaxy of SN 2025bvm, NGC 4156, also fell within the same field of view SN 2025bvm was discovered by the Zwicky Transient Facility (ZTF; Bellm2019) on February 20, 2025, at coordinates RA = 12:10:48.75, DEC = +39:28:28.0 (J2000.0). The Colibri telescope monitored approximately 121 days, from February 23 to June 24, 2025, obtaining photometric data for 79 nights in the $\mathrm{B}$, $\mathrm{g^{\prime}}$, $\mathrm{r^{\prime}}$, and $\mathrm{i^{\prime}}$bands.
@@
All images obtained in these observations were processed using standard procedures, including bias subtraction and flat-field correction. This data reduction was performed using IRAF (tody1986) and Python packages, specifically Astropy (astropy2013; astropy2018; astropy2022). Instrumental magnitudes were extracted using differential photometry, which relies on comparing the brightness of the supernova to a set of reference stars within the same field. Photometric calibration used a set of standard stars from the supernova field, identified in Figure 1 and listed in Table 1. The primary references used were stars BD+40 2507 (marked as 2) and UI+40-42 (marked as 3). The magnitudes for these standard stars in the corresponding filters were obtained from public catalogs such as the Sloan Digital Sky Survey (Abdurro2022) and SIMBAD astronomical database [^1] https://vizier.cds.unistra.fr/viz-bin/VizieR-3?-source=II/336/apass9.
@@
Figure 1 shows the full field of view, indicating the positions of supernova SN 2025bvm (7), its host galaxy NGC 4156 (6), the nearby galaxy NGC 4151 (1), and the reference stars. A detailed log of all photometric observations is presented in Table 2.